The following direct labor standards have been established for a particular product:

Standard labor rate ..........................    $15.00 per hour

Standard labor hours per unit............. 0.5 hours per unit

The following actual data has occurred:

Actual products manufactured.................   4,000 units

Actual hours worked................................. 2,500 hours

Actual total labor cost................................. $35,000

Using a "goalpost" like we did in class, answer the following questions, making sure to indicate if the variance is favorable or unfavorable. You do not have to type in a goalpost, but put your answer for each question (a goalpost on some scrap paper will help you solve for the answer).

a.   What is the total direct labor variance for the month?

b.   What is the direct labor rate variance for the month?

c.   What is the direct labor efficiency variance for the month?
Answer #1

Total labor variance = (SR*AH) - (AR*AH)

= (15*0.5*4000) - 35,000

= 5000 Unfavorable

Labor Rate Variance = (SR-AR) *AH

= (15 - 35,000/2,500)*2500

= 2500 Favourable

Labor efficiency variance = (SH-AH) * SR

= (4,000*0.5 - 2500) * 15

= 7500 Unfavorable